What is the zero to peak amplitude in (N/C) of an electric field if there is an electromagnetic wave in free space with a time average intensity of 60 watts/meter squared?
Answer #1

Average intensity is given by, I = \frac{E^2}{2 \mu_o c} ​​​​​​

Peak amplitude of electric field is, E = \sqrt{2 \mu_o c I} ​​​​​​

\Rightarrow E = \sqrt{2 *4 \pi * 10^{-7}*3*10^8 *60 }=212. 7N/C